OpenStudy (anonymous):

What is equal to 3[f(2x)] for f(x) = 3x^2+3/2

OpenStudy (anonymous):

substitute 2x instead of x and multiply the whole equation so obtained into 3

OpenStudy (anonymous):

f(x)= 3(x)^2 +3/2 f(2x)=3(2x)^2 +3/2=12(x)^2 +3/2 3f(2x)=36(x)^2 +9/2